# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ 2.3 Modelling Features
Broadly, modelling features can be categorized into static and dynamic. For static elements, the actions of traffic agents depend on the physical layout of the environment. Methods that do not account for the environment’s topology assume agents operate in an open-space environment. For the dynamic elements, prediction models extract features related to the target agents and their interactions. Further, we elaborate on each category.
@@ 2.3.2 Target Agents
As discussed in Section 2.2, prediction methods generate forecasts based on the past trajectories of target agents [84, 85, 86, 87], and/or incorporate velocity history [88, 41, 89], as well as heading and orientation [90, 91], to predict future trajectories. In heterogeneous models, including the category of the object (e.g., person, truck, cyclist) enables prediction models [92, 93] to learn category-specific dynamics, although these methods may be prone to errors caused by the incorrect target agent classification by the detector. For vehicles, learning individual behavioral patterns and motion dynamics from historical data [94] can improve prediction accuracy by capturing unique driving styles. Similarly, in pedestrian trajectory prediction, appearance and posture offer valuable insights into pedestrians’ intentions and planned trajectories. Studies [95, 96, 97] have focused on extracting visual cues to capture behavioral patterns and improve prediction accuracy. To estimate posture, two main approaches are employed: using skeleton key-points [50, 98, 99, 100] or leveraging pretrained CNNs [101, 102, 98].
@@ 2.3.3 Interaction Features
This section discusses Computational Cognition and Social Pooling methods, while GNN-based modelling of interactions between traffic agents is addressed in Section 3.4. Social Pooling is a widely adopted learning-based method in the literature. For a comprehensive review of other approaches and a detailed study on interaction modelling for autonomous driving, readers are referred to the most extensive review available on this topic [109].
@@
To implement machine ToM, Inverse Reinforcement Learning (IRL) provides a framework to model how agents’ actions are influenced by the presence of others. Sun et al. [113] applied hierarchical IRL to learn distributions of future trajectories from human demonstrations, capturing multi-agent interactions. Ozkan and Ma [114] introduced a control design for autonomous vehicles that accounts for socially desirable behavior in car-following scenarios, embedding the influence of one agent on another. Tian et al. [115] extended this approach by proposing an IRL framework to infer humans’ latent intelligence levels, demonstrating how agents reason about mutual decision-making in game-theoretic scenarios.
@@ 2.4 Modelling Approach
On more recent approaches, diffusion models [133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143] have been actively utilized in trajectory forecasting applications. These models involve a two-phase process: a forward step where Gaussian noise is progressively added to the data, transforming it into a noisy latent representation aligned with a simple prior distribution, and a reverse step where the model iteratively learns to denoise this representation conditioned on past trajectories and scene context. During inference, random Gaussian noise serves as the starting point, and the denoising process progressively refines it into a plausible future trajectory. By conditioning on contextual information such as past motion and environmental features, diffusion models generate diverse, multimodal, and context-aware trajectory predictions. Jiang et al. [137] proposed MotionDiffuser, which leverages a permutation-invariant transformer-based denoiser with cross-attention mechanisms to integrate scene context, past trajectories, and agent interactions. It employs PCA to compress trajectories into a latent space for computational efficiency and supports controlled sampling via differentiable cost functions, enabling constraints such as collision avoidance. Barquero et al. [138] introduced the BeLFusion model, a diffusion framework for behavior-driven human motion prediction, which uses a behavior coupler to conditionally sample from a behavioral latent space, ensuring realistic and behaviorally coherent trajectories. Yuan et al. [136] developed PhysDiff, integrating a motion projection module with a physics simulator to iteratively guide the denoising process, eliminating artifacts like floating and ground penetration, and producing physically accurate trajectories.
@@ 2.5 Output Modalities
The primary classification of trajectory prediction outputs is based on whether they are unimodal or multimodal. Unimodal outputs are deterministic trajectory points or parameters of a unimodal trajectory distribution. Several studies [22, 92, 27, 24] focus on generating unimodal predictions. However, a significant limitation of unimodal approaches is their inability to capture the inherently multimodal nature of driving behaviour. For instance, in a lane with arrows indicating both straight and left turns, a vehicle could either proceed straight or turn left. Until the driver signals their intent (e.g., by activating an indicator or beginning to incline left), both options should be considered as equally plausible. Ignoring these multimodal characteristics makes it difficult to ensure collision avoidance in complex traffic scenarios, such as lane merging or navigating intersections.
Multimodal outputs address this limitation by generating predictions as a multimodal probability distribution [19, 20, 64]. These outputs better align with real-world traffic scenarios, though they present greater challenges in training. A key advantage of multimodal distributions is their ability to associate probabilities with each possible trajectory. This capability allows models to reflect prediction confidence. For example, in scenarios where an object is partially occluded or just entering the field of view, prediction certainty naturally decreases. For objects, that just entered the scene, robust predictions require past trajectory data of at least 7–10 frames; before this is achieved, models often pad past trajectories with zeros to the desired length. This lack of historical data should be accounted for and reflected in the output probabilities. Furthermore, by explicitly modelling uncertainty, multimodal predictions improve downstream planning by enabling risk-aware decision-making, such as adjusting speed during ambiguous situations or prioritizing safer maneuvers.
@@
A portion of studies focuses on predicting manoeuvre intentions as output [18, 17, 21, 26, 60]. Common manoeuvre categories include changing lanes, staying in the current lane, going straight, and turning left or right. For the joint prediction of trajectories and manoeuvres, two main approaches exist. The first approach involves multi-task models [160], where manoeuvre intentions and trajectories are predicted as separate outputs but share a common feature extraction module. The second approach [161, 162, 163, 101, 102] predicts manoeuvres and trajectories separately, with the predicted manoeuvre used to guide trajectory prediction.
@@
Lastly, occupancy grid and occupancy flow are alternative output formats for prediction models. These methods [164, 165, 166] estimate the future occupancy of each cell in a BEV map of the driving area. Occupancy grids provide a spatial representation by predicting the probability of each grid cell being occupied at future time steps [167, 168] and directly estimating spatial probabilities for the entire scene. However, the accuracy of these predictions is constrained by the resolution of the grid cells. Increasing the number of cells improves precision but significantly raises computational demands. A potential solution is to use a flexible grid resolution that adapts to the environment: finer grids for dense urban traffic and coarser grids for highways, optimizing the balance between precision and computational efficiency. Occupancy flow extends the concept of occupancy grids by not only estimating the future occupancy of grid cells but also capturing the motion dynamics within these cells [169, 170, 171]. While occupancy grids encode only spatial information, occupancy flow incorporates temporal changes by predicting both the future occupancy and the flow, reflecting how the objects will move. By modelling motion dynamics at the grid-cell level, occupancy flow eliminates the need for explicitly tracking individual past trajectories, as it encodes both spatial and temporal information directly from the input data. These properties, make both these formats suited for complex traffic scenarios and smooth integration with planning module.
@@ 2.8 Prediction Paradigms
Prediction models do not exist in isolation; it is an integral part of the autonomous driving stack and must be designed with its role in perception and planning in mind. To ensure feasibility for real-world deployment, prediction models should seamlessly integrate into the broader system architecture. This section discusses three existing prediction paradigms: Detect-Track-Predict, End-to-End Approaches, and Query-Based Approaches, which are graphically illustrated in Figure 6. For each paradigm, we outline the general concept and highlight its advantages and limitations for integration.
@@ 2.8.1 Detect-Track-Predict
The Detect-Track-Predict paradigm is a modular cascade approach consisting of detection, tracking, and prediction modules. First, the detection module localizes objects from sensor data. Next, the tracking module initializes tracklets for the detected objects and maintains their trajectories consistently over time. Finally, the prediction module takes the active tracklets and forecasts their future positions. This approach is one of the most common implementations in autonomous driving and is widely used in the literature. Its modular design makes it easier to debug, as most datasets provide ground truth annotations for each of the modules. Additionally, the extensive research output for detectors [33] and trackers [191], covering both 2D and 3D modalities, further supports its seamless development.
The latest detection approaches leverage data fusion from cameras and LiDAR sensors to enhance detection and tracking accuracy as shown in Figure 7. Early Fusion integrates raw sensor data, such as images and point clouds. A common implementation involves detecting objects in the image, projecting the point cloud onto the 2D image plane, and matching the projected points with the detected objects to classify them and measure distances accurately. Guidolini et al. [192] proposed such a pipeline for combining the strengths of both modalities. In contrast, Late Fusion operates at the detection stage, combining results from LiDAR and camera detections. This can be implemented in two ways: (1) merging 3D bounding boxes detected independently from LiDAR and camera, or (2) projecting LiDAR’s 3D bounding boxes onto the 2D image plane and aligning them with the 2D detections from the camera. This strategy provides flexibility, allowing fusion to occur in either the two-dimensional or three-dimensional space, depending on the accuracy of the detection techniques for both modalities. Another fusion strategy, Augmenting Point Clouds, enhances LiDAR data by projecting point clouds onto images and enriching each point with matching image features. This augmented representation is then fed into a LiDAR-only detection method. Vora et al. [193] proposed PointPainting, which projects LiDAR points onto semantic segmentation outputs from the camera and appends the corresponding class scores to each point. Camera-to-LiDAR projections can result in a loss of rich semantic information. Bird’s-eye view (BEV) Fusion mitigates this issue by combining features from different modalities within a unified BEV representation, providing a more comprehensive spatial understanding. For a detailed overview, we refer the reader to the multi-sensor fusion framework presented in [194].
@@
Kalman Filter-based [195, 196, 197], leverage predicted object states to mitigate detection inaccuracies caused by sensor noise or missed detections. Advanced trackers [198] address partial occlusions, ensuring robust trajectory maintenance and minimizing false positives that lead to ghost tracklets. Prediction models rely on the precision of trajectories over the past 10–15 frames, making robust tracklets essential for accurate forecasting. Excessive ghost tracklets result in non-existent predictions being passed to the planner, potentially degrading overall system performance.
@@ 2.8.2 End-to-End Models
By sharing high-dimensional scene features and utilizing components like the pool, end-to-end models retain richer contextual information, enabling the learning of intermediate representations that benefit both detection and prediction tasks. This paradigm avoids the information bottlenecks of cascade approaches by eliminating the tracking module and jointly optimizing detection and prediction. However, despite these advantages, errors in the shared features from the detection stage can propagate through the system, affecting overall performance.
@@ 2.8.3 Query-Based Prediction Models
Query-Based prediction models represent a cutting-edge approach to trajectory prediction by leveraging learnable queries within a unified framework. Rather than splitting detection and prediction into separate stages, query-based models directly focus on objects of interest, iteratively refining their trajectory predictions. These learnable object queries encapsulate spatial and temporal parameters, guiding the model in producing accurate and consistent predictions. By integrating contextual data over multiple refinement steps, query-based models capture fine-grained details and interaction dynamics. A key advantage of this approach is the elimination of error propagation across separate subsystems, as predictions are generated holistically within a single framework. In the following discussion, we highlight several query-based methods proposed in the literature.
@@ 3 Survey of Modelling Methods
This section describes the proposed taxonomy of prediction approaches. To systematically categorize a broad range of prediction models from the literature, we structure our taxonomy based on the modeling approach, providing a concise classification and clustering of methods. Selecting an appropriate modeling approach based on application requirements and motion modeling aspects is crucial for designing an efficient, robust, and competitive prediction model comparable to SOTA approaches.
@@ 3.1 Taxonomy
In the proposed taxonomy, prediction methods are divided based on the way motion is modeled, i.e., physics-based methods explicitly define motion equations, while learning-based approaches learn motion patterns from observations. Learning-based approaches are further subdivided into classical machine learning, deep learning, and reinforcement learning-based methods (please refer to Figure 8). Each category of methods is elaborated upon below.
@@
    Physics-based methods predict motion by simulating a set of equations describing a physics-based model. Single-Trajectory methods predict trajectory by directly applying dynamic or kinematic equations to the agent’s current state. Kalman Filtering methods account for the agent’s state uncertainty, which increases during prediction steps and reduces after receiving observations. Monte Carlo methods approximate the state distribution by randomly sampling from the input variables and applying a physics model to generate potential future trajectories. Multiple Model methods account for motion-mode uncertainty and combine several motion models to better describe the complexity of the agent’s motion behavior.
@@
    Learning-based methods, in contrast to physics-based approaches, do not exploit explicitly defined functions governing the motion but instead learn motion dynamics from driving data.
@@
      Deep Learning (DL) methods also learn motion dynamics from driving data but are more suitable for long-term predictions in complex environments than conventional ML models, due to their capability of learning complex and hierarchical data representations. Among DL methods, the following main groups of models are reviewed: Sequential Models, Convolutional Neural Networks, Graph Neural Networks, Transformers, AutoEncoders, and Generative Adversarial Networks.
@@
      Reinforcement Learning (RL) methods interpret driving as a high-dimensional complex policy. Despite these methods having direct applications for planning, recent works increasingly utilize RL methods for trajectory prediction through learning via a reward system. Inverse Reinforcement Learning methods learn a reward function based on expert demonstrations. Imitation Learning bypasses the stage of learning reward functions and directly extracts an optimal policy from the data, generating future trajectories similar to expert trajectories.
Figure 8: Overview of the Taxonomy of modelling methods.

### v2

@@ 2.3 Modelling Features
Modeling features in trajectory prediction can be broadly categorized into static and dynamic components. Static features refer to the physical layout of the environment, such as road geometry, sidewalks, or walkable areas. Models that include this information can incorporate spatial constraints, e.g., road boundaries, lanes, or walkable regions, into the forecasting process, allowing predicted trajectories to better align with the underlying scene layout. Dynamic features include features related to the target agents and their interactions. One such feature is the agent’s past motion, where the number of past steps is a design choice and varies across models. For example, physics-based models typically rely on the last one or two steps, while deep learning models with attention mechanisms can learn temporal dependencies from longer sequences. Interaction modeling can be beneficial in the scenarios of dense traffic and intersections, where agent behaviors are tightly coupled. In the following sections, we elaborate on each category and describe how different models employ these features.
@@ 2.3.2 Target Agents
As discussed in Section 2.2, it is common for prediction methods to generate forecasts based on the past trajectories of target agents [84, 85, 86, 87], and/or incorporate velocity history [88, 41, 89], as well as heading and orientation [90, 91]. In heterogeneous models, including the category of the object (e.g., person, truck, cyclist) enables prediction models [92, 93] to learn category-specific dynamics, although these methods may be prone to errors caused by the incorrect target agent classification by the detector. For vehicles, learning individual behavioral patterns and motion dynamics from historical data [94] can improve prediction accuracy by capturing unique driving styles. Similarly, in pedestrian trajectory prediction, appearance and posture can offer insights into pedestrians’ intentions and planned trajectories. Studies [95, 96, 97] have focused on extracting visual cues to capture behavioral patterns and improve prediction accuracy. To estimate posture, two main approaches are employed: using skeleton key-points [50, 98, 99, 100] or leveraging pretrained CNNs [101, 102, 98].
@@ 2.3.3 Interaction Features
This section briefly elaborates on Computational Cognition and Social Pooling methods, while GNN-based modelling of interactions between traffic agents is addressed in Section 3.4. Social Pooling is a widely adopted learning-based method in the literature. For a comprehensive review of other approaches and a detailed study on interaction modelling for autonomous driving, readers are referred to the most extensive review available on this topic [109].
@@ 2.4 Modelling Approach
On more recent approaches, diffusion models [130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140] have been actively utilized in trajectory forecasting. These models follow a two-phase process: first, Gaussian noise is added to the data to produce a latent representation aligned with a simple prior; then, a reverse process iteratively denoises it, conditioned on past trajectories and scene context. During inference, starting from random noise, the model refines predictions into diverse, multimodal, and context-aware trajectories. Jiang et al. [134] proposed MotionDiffuser, which leverages a permutation-invariant transformer-based denoiser with cross-attention mechanisms to integrate scene context, past trajectories, and agent interactions. It employs PCA to compress trajectories into a latent space for computational efficiency and supports controlled sampling via differentiable cost functions, enabling constraints such as collision avoidance. Barquero et al. [135] introduced the BeLFusion model, a diffusion framework for behavior-driven human motion prediction, which uses a behavior coupler to conditionally sample from a behavioral latent space, ensuring realistic and behaviorally coherent trajectories. Yuan et al. [133] developed PhysDiff, integrating a motion projection module with a physics simulator to iteratively guide the denoising process, eliminating artifacts like floating and ground penetration, and producing physically accurate trajectories.
@@ 2.5 Output Modalities
The primary classification of trajectory prediction outputs is based on whether they are unimodal or multimodal. Unimodal outputs are deterministic trajectory points or parameters of a unimodal trajectory distribution. Several studies [22, 92, 27, 24] focus on generating unimodal predictions. However, a significant limitation of unimodal approaches is their inability to capture the inherently multimodal nature of driving behaviour. For instance, in a lane with arrows indicating both straight and left turns, a vehicle could either proceed straight or turn left. Until the driver signals their intent (e.g., by activating an indicator or beginning to incline left), both options should be considered as equally plausible. Ignoring these multimodal characteristics makes it difficult to ensure collision avoidance in complex traffic scenarios, such as lane merging or navigating intersections. Multimodal outputs address this limitation by generating predictions as a multimodal probability distribution [19, 20, 64]. These outputs better align with real-world traffic scenarios, though they present greater challenges in training. A key advantage of multimodal distributions is their ability to associate probabilities with each possible trajectory.
@@
A portion of studies focuses on predicting maneuver intentions as output [18, 17, 21, 26, 60]. Common maneuver categories include changing lanes, staying in the current lane, going straight, and turning left or right. For the joint prediction of trajectories and maneuvers, two main approaches exist. The first approach involves multi-task models [157], where maneuvers intentions and trajectories are predicted as separate outputs but share a common feature extraction module. The second approach [158, 159, 160, 101, 102] predicts maneuvers and trajectories separately, with the predicted maneuvers used to guide trajectory prediction.
@@
Lastly, occupancy grid and occupancy flow are alternative output formats for prediction models. These methods [161, 162, 163] estimate the future occupancy of each cell in a BEV map of the driving area. Occupancy grids provide a spatial representation by predicting the probability of each grid cell being occupied at future time steps [164, 165] and directly estimating spatial probabilities for the entire scene. However, the accuracy of these predictions is constrained by the resolution of the grid cells. A potential solution is to use a flexible grid resolution that adapts to the environment: finer grids for dense urban traffic and coarser grids for highways, optimizing the balance between precision and computational efficiency. Occupancy flow extends the concept of occupancy grids by not only estimating the future occupancy of grid cells but also capturing the motion dynamics within these cells [166, 167, 168]. While occupancy grids encode only spatial information, occupancy flow incorporates temporal changes by predicting both the future occupancy and the flow, reflecting how the objects will move. By modelling motion dynamics at the grid-cell level, occupancy flow eliminates the need for tracker, as it encodes both spatial and temporal information directly from the input data. Both formats are well suited for complex traffic scenarios and smooth integration with planning module.
@@ 2.8 Prediction Paradigms
This section discusses three existing prediction paradigms: Detect-Track-Predict, End-to-End Approaches, and Query-Based Approaches, which are graphically illustrated in Figure 6. For each paradigm, we outline the general concept and highlight its advantages and limitations for integration.
@@ 2.8.1 Detect-Track-Predict
The Detect-Track-Predict paradigm is a modular cascade approach consisting of detection, tracking, and prediction modules. First, the detection module localizes objects from sensor data. Next, the tracking module initializes tracklets for the detected objects and maintains their trajectories consistently over time. Finally, the prediction module takes the active tracklets and forecasts their future positions. This approach is one of the most common implementations in autonomous driving and is broadly used in the literature. Its modular design makes it easier to debug, as most datasets provide ground truth annotations for each of the modules. Additionally, the extensive research output for detectors [33] and trackers [188], covering both 2D and 3D modalities, further supports its seamless development.
The latest detection approaches leverage data fusion from cameras and LiDAR sensors to enhance detection and tracking accuracy as shown in Figure 7. Early Fusion integrates raw sensor data, such as images and point clouds. A common implementation involves detecting objects in the image, projecting the point cloud onto the 2D image plane, and matching the projected points with the detected objects to classify them and measure distances accurately. Guidolini et al. [189] proposed such a pipeline for combining the strengths of both modalities. In contrast, Late Fusion operates at the detection stage, combining results from LiDAR and camera detections. This can be implemented in two ways: (1) merging 3D bounding boxes detected independently from LiDAR and camera, or (2) projecting LiDAR’s 3D bounding boxes onto the 2D image plane and aligning them with the 2D detections from the camera. This strategy provides flexibility, allowing fusion to occur in either the two-dimensional or three-dimensional space, depending on the accuracy of the detection techniques for both modalities. Another fusion strategy, Augmenting Point Clouds, enhances LiDAR data by projecting point clouds onto images and enriching each point with matching image features. This augmented representation is then fed into a LiDAR-only detection method. Vora et al. [190] proposed PointPainting, which projects LiDAR points onto semantic segmentation outputs from the camera and appends the corresponding class scores to each point. Camera-to-LiDAR projections can result in a loss of rich semantic information. Bird’s-eye view (BEV) Fusion mitigates this issue by combining features from different modalities within a unified BEV representation, providing a more comprehensive spatial understanding, but resulting in a prolonged inference time. For a detailed overview, we refer the reader to the multi-sensor fusion framework presented in [191].
@@
Kalman Filter-based trackers [192, 193, 194], leverage predicted object states to mitigate detection inaccuracies caused by sensor noise or missed detections. Advanced trackers [195] address partial occlusions, ensuring robust trajectory maintenance and minimizing false positives that lead to ghost tracklets. Prediction models rely on the precision of trajectories over the past 10–20 timestamps, making robust tracklets essential for accurate forecasting. Excessive ghost tracklets result in non-existent predictions being passed to the planner, potentially degrading overall system performance.
@@ 2.8.2 End-to-End Models
By sharing high-dimensional scene features, end-to-end models retain richer contextual information, enabling the learning of intermediate representations that benefit both detection and prediction tasks. This paradigm avoids the information bottlenecks of cascade approaches by eliminating the tracking module and jointly optimizing detection and prediction. However, despite these advantages, errors in the shared features from the detection stage can propagate through the system, affecting overall performance.
@@ 2.8.3 Query-Based Prediction Models
Query-Based prediction models represent a cutting-edge approach to trajectory prediction by leveraging learnable queries within a unified framework. Rather than splitting detection and prediction into separate stages, query-based models directly focus on objects of interest, iteratively refining their trajectory predictions. These learnable object queries encapsulate spatial and temporal parameters, guiding the model in producing consistent predictions. By integrating contextual data over multiple refinement steps, query-based models capture fine-grained details and interaction dynamics. A key advantage of this approach is the elimination of error propagation across separate subsystems, as predictions are generated holistically within a single framework. In the following discussion, we highlight several query-based methods proposed in the literature.
@@ 3 Survey of Modelling Methods
This section describes the refined taxonomy of trajectory prediction approaches. To systematically categorize the wide range of prediction models in the literature, the taxonomy is structured based on modeling methods. It builds upon the categorizations presented in reviews [4, 5], and is further extended based on a review of recent works in which prediction models were systematically tagged and grouped.
@@ 3.1 Taxonomy
In the proposed taxonomy, prediction methods are divided based on the way motion is modeled, i.e., physics-based methods explicitly define motion equations, while learning-based approaches learn motion patterns from observations. Learning-based approaches are further subdivided into classical machine learning, deep learning, and reinforcement learning-based methods (please refer to Figure 8).
@@
    Physics-based methods predict motion by simulating a set of equations describing a physics-based model. Single-Trajectory methods predict trajectory by directly applying dynamic or kinematic equations to the agent’s current state. Kalman Filtering methods account for the agent’s state uncertainty, which increases during prediction steps and reduces after receiving observations. Monte Carlo methods approximate the state distribution by randomly sampling from the input variables and applying a physics model to generate potential future trajectories. Multiple Model methods account for motion-mode uncertainty and combine several motion models to better describe the complexity of agent motion behavior.
@@
    Learning-based methods, in contrast to physics-based approaches learn motion dynamics from driving data.
@@
      Deep Learning (DL) methods also learn motion dynamics from driving data by leveraging their ability to capture complex and hierarchical patterns. Among DL methods, the following main groups of models are reviewed: Sequential Models, Convolutional Neural Networks, Graph Neural Networks, Transformers, AutoEncoders, and Generative Adversarial Networks.
@@
      Reinforcement Learning (RL) methods interpret driving as a high-dimensional complex policy. Despite these methods having direct applications for planning, recent works attempt to utilize RL methods for trajectory prediction through learning via a reward system. Inverse Reinforcement Learning methods learn a reward function based on expert demonstrations. Imitation Learning bypasses the stage of learning reward functions and directly extracts an optimal policy from the data, generating future trajectories similar to expert trajectories.
Figure 8: Taxonomy of modeling methods for trajectory prediction, divided into two main branches: physics-based approaches and learning-based approaches. The learning-based category is further subdivided into classical machine learning methods, deep learning architectures and reinforcement learning approaches.